# Stable and Accurate Orbital-Free DFT Powered by Machine Learning

### v1

@@ Acknowledgments
The authors would like to thank the STRUCTURES excellence cluster for nurturing a unique collaborative environment that made this work possible in the first place. The authors would also like to thank Maurits Haverkort for helpful discussions, and Corinna Steffen for help with creating plots and for chemical insight.
@@
This work is supported by Deutsche Forschungsgemeinschaft (DFG) under Germany’s Excellence Strategy EXC-2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster). The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG. C.A.G. and M.V.K. were supported by the Konrad Zuse School of Excellence in Learning and Intelligent Systems (ELIZA) through the DAAD program Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research. T.K. and A.D. acknowledge support by the German Research Foundation (DFG) through grant no INST40/575-1 FUGG (JUSTUS 2 cluster). The authors also thank SFB 1249 for partial funding as well as the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) for Projektnummer 240245660 - SFB 1129. We also thank Klaus Tschira Stiftung gGmbH for support in the framework of the SIMPLAIX consortium.
@@
Appendix A Materials and Methods
A.1 Data generation details
A.1.1 Datasets
@@
The extrapolation capabilities of STRUCTURES25 are tested on the QMugs dataset [38]. The latter contains more than 665k drug-like molecules from the ChEMBL database. We filter out sulfur, chlorine, bromine and iodine since these elements do not appear in the QM9 dataset. For comparison with [13], we split the dataset into multiple bins according to size. The first bin comprises molecules with 10 to 15 heavy atoms and is used for training. The following bins have a width of 5 heavy atoms and contain 50 randomly sampled molecules each which are used in density optimization.
A.1.2 Kohn-Sham DFT settings
For the training and test label generation, KS-DFT calculations were carried out using the open source software package PySCF [47, 48, 49]. For ease of comparison with prior work M-OFDFT [13], we largely choose identical hyperparameters. Restricted-spin calculations employing the 6-31G(2df,p) basis set [50] were conducted using the established general gradient approximation (GGA) functional PBE [51] with a grid level of 3 for QM9 and a grid level of 2 for QMugs, respectively. For larger molecules with more than 30 atoms, density fitting with the def2-universal-jfit basis set was enabled. We set the convergence tolerance to the PySCF default of $2.72\cdot 10^{-5}$ meV for QM9 and to 1 meV for QMugs. We use the commutator direct inversion of the iterative subspace (C-DIIS)  [52, 53] method with a maximal subspace of eight iterations. Minimal atomic orbitals (MINAO) was used as initialization [43, 44]. The open source nature and python implementation of PySCF enabled us to insert callbacks in between the SCF steps of the KS-DFT computation. These serve two purposes: First, to extract density labels, energy labels, gradient labels and DIIS coefficients. Secondly and implicitly, to add perturbations to the Fock matrix, slowing down the convergence and leading to more varied training labels (see section A.1.3).
@@ A.1.2 Kohn-Sham DFT settings
$\left(-\frac{1}{2}\nabla^{2}+V_{\mathrm{eff}}\left[\{\phi_{i}\}\right]\right)\phi_{i}=\varepsilon_{i}\phi_{i}$ (A.1)
@@
$\mathbf{F}^{t}\mathbf{C}^{t}=\mathbf{S}\mathbf{C}^{t}\boldsymbol{\varepsilon}^{t},\qquad F_{\alpha\beta}^{t}=\int\mathrm{d}^{3}\mathbf{r}\,\eta_{\alpha}(\mathbf{r})\left[-\frac{1}{2}\nabla^{2}+V_{\mathrm{eff}}[\{\phi^{t-1}_{i}\}](\mathbf{r})\right]\eta_{\beta}(\mathbf{r}),$ (A.2)
@@
$\rho_{\mathrm{OF}}(\mathbf{r})=\sum_{\mu=1}^{N_{\mathrm{OF}}}p_{\mu}\omega_{\mu}(\mathbf{r}),$ (A.3)
where $p_{\mu}$ are new coefficients that can be obtained from $C_{\alpha i}$ using density fitting as described in section A.1.4.
A.1.3 Perturbation of the effective potential
@@
Our main contribution is to instead perturb the effective potential $V_{\mathrm{eff}}$, which is used in each SCF iteration to generate the electron density of the next iteration (cf.  Eqs. A.2 and A.4). This approach is simple and stable, and results in labeled data that is much more evenly distributed than densities generated naively from Kohn-Sham DFT. It also offers direct control of the strength of perturbation in the effective potential $V_{\mathrm{eff}}$ which in turn correlates with the strength of perturbation in electron densities and energies. This is illustrated in Fig. 3C, where the difference between the non-interacting kinetic energy $T_{\mathrm{S}}$ and the corresponding value at the ground state $T_{\mathrm{S}}^{\mathrm{gs}}$ is shown, using both perturbed and unperturbed $V_{\mathrm{eff}}$. In both cases, samples are generated from all molecular geometries in the QM9 validation set, the higher total number of samples in the former case stems from the increased number of SCF iterations per molecule due to the perturbations.
The effective potential $V_{\mathrm{eff}}$ in Eq. A.1 is perturbed from the sixth up to the 26th SCF iteration, counting the initial guess as the zeroth iteration. Only these perturbed samples are used in training. The perturbed Kohn-Sham equations read $\left[-\frac{1}{2}\nabla^{2}+V_{\mathrm{eff}}[\{\phi_{i}^{t-1}\}]+\Delta^{t}\right]\phi^{t}_{i}=\varepsilon^{t}_{i}\phi^{t}_{i}$, with the perturbation function $\Delta(\mathbf{r})=\sum_{\mu}d_{\mu}\omega_{\mu}(\mathbf{r})$, its coefficients $d_{\mu}$ are sampled from a normal distribution with a standard deviation decreasing linearly from $0.102$ to $0.002$ over the SCF iterations and then multiplied with the orbital-free basis functions $\{\omega_{\mu}\}$. In the $\{\eta_{\alpha}\}$ basis representation this amounts to
$\left(\mathbf{F}^{t}+\mathbf{\Delta}^{t}\right)\mathbf{C}^{t}$ $=\mathbf{S}\mathbf{C}^{t}\boldsymbol{\varepsilon}^{t}$ (A.4)
$\Delta^{t}_{\alpha\beta}$ $=\sum_{\mu}d^{t}_{\mu}\int\mathrm{d}^{3}r\,\eta_{\alpha}(\mathbf{r})\eta_{\beta}(\mathbf{r})\omega_{\mu}(\mathbf{r}).$ (A.5)
@@ A.1.4 Label generation
To obtain density coefficients in the orbital-free ansatz (Eq. 2), we employ density fitting. The resolution of identity method by Whitten[54] and Dunlap[55] can be used to fit the orbital-free density coefficients $\mathbf{p}=\{p_{\mu}\}_{\mu\in 1,\dots,N_{\mathrm{OF}}}$ to a Kohn Sham density. Here, the discrepancy of the fit is measured by the residual Hartree energy
$E_{\mathrm{H}}[\rho_{\mathrm{KS}}-\rho_{\mathrm{OF}}]$ $=\int\mathrm{d}\mathbf{r}\int\mathrm{d}\mathbf{r^{\prime}}\frac{(\rho_{\mathrm{KS}}(\mathbf{r})-\rho_{\mathrm{OF}}(\mathbf{r}))(\rho_{\mathrm{KS}}(\mathbf{r^{\prime}})-\rho_{\mathrm{OF}}(\mathbf{r^{\prime}}))}{|\mathbf{r}-\mathbf{r^{\prime}}|}$ (A.6)
$=\mathbf{p}^{\top}\mathbf{\tilde{W}}\mathbf{p}-2\mathbf{p}\,\mathrm{tr}(\mathbf{\tilde{L}}\mathbf{\Gamma})+\sum\limits_{\alpha\beta\gamma\delta}\Gamma_{\alpha\beta}\tilde{D}_{\alpha\beta,\gamma\delta}\Gamma_{\gamma\delta}.$ (A.7)
@@
$(f|g)=\int\mathrm{d}\mathbf{r}\int\mathrm{d}\mathbf{r}^{\prime}\frac{f(\mathbf{r})g(\mathbf{r^{\prime}})}{|\mathbf{r}-\mathbf{r^{\prime}}|}.$ (A.8)
@@
As Eq. A.7 is a quadratic form in $\mathbf{p}$, it can be minimized analytically. But in agreement with M-OFDFT[13], we found that also considering the residual external energy in the minimization yields better fitted external and exchange correlation energies, as using the residual Hartree energy alone only leads to a close fit for this energy. The residual external energy reads
$E_{\mathrm{ext}}[\rho_{\mathrm{KS}}-\rho_{\mathrm{OF}}]$ $=\int\mathrm{d}\mathbf{r}\sum\limits_{(Z_{a},\mathbf{R}_{a})\in\mathcal{M}}\frac{-Z_{a}(\rho_{\mathrm{KS}}(\mathbf{r})-\rho_{\mathrm{OF}}(\mathbf{r}))}{|\mathbf{R}_{a}-\mathbf{r}|}$ (A.9)
$=\mathbf{v}_{\mathrm{ext}}^{\top}\,\mathbf{p}-\mathrm{tr}(\Gamma\mathbf{V}_{\mathrm{ext}}),$ (A.10)
with $[{\mathbf{v}_{\mathrm{ext}}}]_{\mu}=E_{\mathrm{ext}}[\omega_{\mu}]$ and $[{\mathbf{V}_{\mathrm{ext}}}]_{\alpha\beta}=E_{\mathrm{ext}}[\eta_{\alpha}\cdot\eta_{\beta}]$ being the external energies of the individual basis functions and $\mathcal{M}$ denoting the molecule geometry, comprising nuclear positions $\mathbf{R}_{a}$ and corresponding charges $Z_{a}$. The objectives of $E_{\mathrm{ext}}[\rho_{\mathrm{KS}}-\rho_{\mathrm{OF}}]=0$ and minimizing $E_{\mathrm{H}}[\rho_{\mathrm{KS}}-\rho_{\mathrm{OF}}]$ can be combined by differentiating Eq. A.7 with respect to $\mathbf{p}$ and coupling with Eq. A.10, yielding an overdetermined linear system111 The M-OFDFT supplementary[13] suggests minimizing the loss $\mathcal{L}(\mathbf{p})=E_{H}[\rho_{\mathrm{OF}}(\mathbf{p})-\rho_{\mathrm{KS}}]+\left(E_{\mathrm{ext}}[\rho_{\mathrm{OF}}(\mathbf{p})]-E_{\mathrm{ext}}[\rho_{\mathrm{KS}}]\right)^{2},$ (A.11) but the available code optimizes the least squares problem as we do. which is solved by minimizing
$\mathcal{L}(\mathbf{p})=\left\lVert\begin{pmatrix}\mathbf{\tilde{W}}\\ \mathbf{v}_{\mathrm{ext}}^{\top}\end{pmatrix}\mathbf{p}-\begin{pmatrix}\mathrm{tr}(\mathbf{\tilde{L}}\mathbf{\Gamma})\\ \mathrm{tr}(\mathbf{\Gamma}\mathbf{V}_{\mathrm{ext}})\end{pmatrix}\right\lVert^{2}$ (A.12)
@@
$E_{\mathrm{tot}}=T_{\mathrm{S}}+E_{\mathrm{XC}}+E_{\mathrm{H}}+E_{\mathrm{ext}}+E_{\mathrm{nuc}},$ (A.13)
@@
$T_{\mathrm{S}}(\mathbf{C})=\sum\limits_{i}\langle\phi_{i}|\hat{T}|\phi_{i}\rangle=\sum_{\alpha,\beta}\sum_{i}C_{\alpha i}C_{\beta i}\langle\eta_{\alpha}|\hat{T}|\eta_{\beta}\rangle.$ (A.14)
As in [13], we calculate the kinetic energy label for the orbital-free density such that the total energy remains constant, $E_{\mathrm{tot}}(\mathbf{p})=E_{\mathrm{tot}}(\mathbf{C})$, yielding
$T_{\mathrm{S}}(\mathbf{p})=T_{\mathrm{S}}(\mathbf{C})+E_{\mathrm{eff}}(\mathbf{C})-E_{\mathrm{eff}}(\mathbf{p}),\qquad E_{\mathrm{eff}}\coloneq E_{\mathrm{H}}+E_{\mathrm{XC}}+E_{\mathrm{ext}},$ (A.15)
@@
$\frac{\delta T_{\mathrm{S}}[\rho]}{\delta\rho(\mathbf{r})}=-V_{\mathrm{eff}}(\mathbf{r})-\Delta(\mathbf{r})+\mu,$ (A.17)
@@
$\nabla_{p_{\nu}}T_{\mathrm{S}}(\mathbf{p})=\int\frac{\delta T_{\mathrm{S}}[\rho]}{\delta\rho(\mathbf{r})}\omega_{\nu}(\mathbf{r})\mathrm{d}\mathbf{r}=\int\left(-V_{\mathrm{eff}}(\mathbf{r})-\Delta(\mathbf{r})+\mu\right)\omega_{\nu}(\mathbf{r})\mathrm{d}\mathbf{r}.$ (A.18)
The unknown chemical potential $\mu$ can be set to zero, as this only yields a gradient contribution orthogonal to the manifold of normalized densities, which is projected out in density optimization (see section A.6). Instead of obtaining the effective potential function $V_{\mathrm{eff}}(\mathbf{r})$ from the coefficients in the orbital basis, we follow M-OFDFT [13] and directly use the effective potential vector given by
$\mathbf{v}_{\mathrm{eff}}(\mathbf{p})=\nabla_{\mathbf{p}}\left(E_{\mathrm{H}}(\mathbf{p})+E_{\mathrm{XC}}(\mathbf{p})+E_{\mathrm{ext}}(\mathbf{p})\right).$ (A.19)
@@
$\nabla_{\mathbf{p}}T_{\mathrm{S}}(\mathbf{p})=-{\mathbf{v}}_{\mathrm{eff}}-\mathbf{W}\mathbf{d},$ (A.20)
@@
$\mathbf{\tilde{F}}^{t}=\sum_{\tau=1}^{t}\pi_{\tau}^{t}\mathbf{F}^{\tau},\qquad\sum_{\tau=1}^{t}\pi_{\tau}^{t}=1.$ (A.21)
@@
$\mathbf{\tilde{V}}_{\mathrm{eff}}^{t}=\sum_{\tau=1}^{t}\pi_{\tau}^{t}\mathbf{V}^{\tau}_{\mathrm{eff}},$ (A.22)
@@
$\tilde{\mathbf{v}}^{t}_{\mathrm{eff}}=\sum_{\tau=1}^{t}\pi_{\tau}^{t}\mathbf{v}^{\tau}_{\mathrm{eff}}.$ (A.23)
@@
A.2 Model overview
The model input consists of a molecular graph with atomic positions $\{\mathbf{R}_{a}\}$ and atom types $\{Z_{a}\}$ as well as the coefficients $\mathbf{p}$, representing the electron density in terms of atom-centered basis functions, cf. Eq. 2. As is customary in local canonicalization, we compute an equivariant local coordinate system for each atom based on the relative position of adjacent non-hydrogen atoms. The basis functions $\{\omega_{\mu}\}$ are a product of a radial function and spherical harmonics. As a consequence, the density coefficients transform via Wigner-D matrices under 3D rotations. To achieve invariance w.r.t. the global orientation of the molecule, the coefficients are transformed into the respective local frame at each atom. As in [13], the coefficients undergo a “natural reparametrization” into an orthonormal basis, that is specific to the molecule geometry (see section A.3 for details); and the coefficients are rescaled dimensionwise to standardize the model input and the desired gradient range of the model w.r.t. the input coefficients.
@@ A.2 Model overview
$\text{ShrinkGate}(\tilde{\mathbf{p}})=\lambda_{\mathrm{out}}\tanh(\lambda_{\mathrm{in}}\,\tilde{\mathbf{p}}),$ (A.24)
@@
$\tilde{e}_{ij}=\eta_{\mathrm{mul}}(Z_{i},Z_{j})\|\mathbf{r}_{i}-\mathbf{r}_{j}\|+\eta_{\mathrm{bias}}(Z_{i},Z_{j})\,,$ (A.25)
$e_{ij}^{k}=\frac{1}{\sqrt{2\pi}\sigma_{k}}\exp\left(\frac{(\tilde{e}_{ij}-\mu_{k})^{2}}{2\sigma_{k}^{2}}\right).$ (A.26)
@@
Further, an embedding of the atom number $Z$ and an aggregation of edge features over neighboring nodes $\operatorname{MLP}\left(\sum_{j\in\mathcal{N}(i)}e_{ij}^{k}\right)$ are added to the node features. These are then passed to a message-passing graph neural network. At the core of the architecture, we apply 4 (or 8 for QMugs) Graphormer blocks [33]. For experiments with local architectures, we propagate messages along a graph with a radial cutoff of 6 Bohr and otherwise use the fully connected graph. Before each attention block, we also apply a node-wise layer norm. Finally, we employ an energy MLP which produces an atom-specific energy contribution from the final node features. Combined with an atom-specific contribution (“atomic reference module”, see below) based on the statistics of the data, the individual energies of each molecule are aggregated by summation into the total energy prediction. For the QMugs model, we used a hierarchical energy readout (similar to [59]), where, instead of a single energy MLP after the final layer, we employ an energy MLP after every second transformer block to predict atom-specific energy contributions. In the end, all energy contributions of the individual readouts are summed into the final atom-wise energies. Having readout MLPs also in intermediate layers, where the field of view is still small, enforces the prediction of more local energy contributions whereas MLPs after later layers are expected to capture increasingly non-local effects.

### v2

@@ Acknowledgments
The authors thank the STRUCTURES excellence cluster for nurturing a unique collaborative environment that made this work possible in the first place. The authors also thank Maurits Haverkort, Corinna Steffen, Virginia Lenk and Andreas Hermann for helpful discussions, chemical insight and analytics.
@@
[…]
@@
S.1 Data generation details
S.1.1 Datasets
@@
The extrapolation capabilities of STRUCTURES25 are tested on the QMugs dataset [49]. The latter contains more than 665k drug-like molecules from the ChEMBL database. We filter out sulfur, chlorine, bromine and iodine since these elements do not appear in the QM9 dataset. For comparison with M-OFDFT [26], we split the dataset into multiple bins according to size. The first bin comprises molecules with 10 to 15 heavy atoms and is used for training. The following bins have a width of 5 heavy atoms and contain 50 randomly sampled molecules each which are used in density optimization.
S.1.2 Kohn-Sham DFT settings
For the training and test label generation, KS-DFT calculations were carried out using the open source software package PySCF [62, 63, 64]. For ease of comparison with prior work M-OFDFT [26], we largely choose identical hyperparameters. Restricted-spin calculations employing the 6-31G(2df,p) basis set [65] were conducted using the established general gradient approximation (GGA) functional PBE [66] with a grid level of 3 for QM9 and a grid level of 2 for QMugs, respectively. For larger molecules with more than 30 atoms, density fitting with the def2-universal-jfit basis set was enabled. We set the convergence tolerance to the PySCF default of $2.72\cdot 10^{-5}$ meV for QM9 and to 1 meV for QMugs. We use the commutator direct inversion of the iterative subspace (C-DIIS)  [67, 68] method with a maximal subspace of eight iterations. Minimal atomic orbitals (MINAO) was used as initialization [55, 56]. The open source nature and python implementation of PySCF enabled us to insert callbacks in between the SCF steps of the KS-DFT computation. These serve two purposes: First, to extract density labels, energy labels, gradient labels and DIIS coefficients. Secondly and implicitly, to add perturbations to the Fock matrix, slowing down the convergence and leading to more varied training labels (see section S.1.3).
@@
$\left(-\frac{1}{2}\nabla^{2}+V_{\mathrm{eff}}\left[\{\phi_{i}\}\right]\right)\phi_{i}=\varepsilon_{i}\phi_{i}$ (S.1)
@@
$\mathbf{F}^{t}\mathbf{C}^{t}=\mathbf{S}\mathbf{C}^{t}\boldsymbol{\varepsilon}^{t},\qquad F_{\alpha\beta}^{t}=\int\mathrm{d}^{3}\mathbf{r}\,\eta_{\alpha}(\mathbf{r})\left[-\frac{1}{2}\nabla^{2}+V_{\mathrm{eff}}[\{\phi^{t-1}_{i}\}](\mathbf{r})\right]\eta_{\beta}(\mathbf{r}),$ (S.2)
@@
$\rho_{\mathrm{OF}}(\mathbf{r})=\sum_{\mu=1}^{N_{\mathrm{OF}}}p_{\mu}\omega_{\mu}(\mathbf{r}),$ (S.3)
where $p_{\mu}$ are new coefficients that can be obtained from $C_{\alpha i}$ using density fitting as described in section S.1.4.
S.1.3 Perturbation of the effective potential
@@
Our main contribution is to instead perturb the effective potential $V_{\mathrm{eff}}$, which is used in each SCF iteration to generate the electron density of the next iteration (cf.  Eqs. S.2 and S.4). This approach is simple and stable, and results in labeled data that is much more evenly distributed than densities generated naively from Kohn-Sham DFT. It also offers direct control of the strength of perturbation in the effective potential $V_{\mathrm{eff}}$ which in turn correlates with the strength of perturbation in electron densities and energies. This is illustrated in Fig. 3C, where the difference between the non-interacting kinetic energy $T_{\mathrm{S}}$ and the corresponding value at the ground state $T_{\mathrm{S}}^{\mathrm{gs}}$ is shown, using both perturbed and unperturbed $V_{\mathrm{eff}}$. In both cases, samples are generated from all molecular geometries in the QM9 validation set, the higher total number of samples in the former case stems from the increased number of SCF iterations per molecule due to the perturbations.
The effective potential $V_{\mathrm{eff}}$ in Eq. S.1 is perturbed from the sixth up to the 26th SCF iteration, counting the initial guess as the zeroth iteration. Only these perturbed samples are used in training. The perturbed Kohn-Sham equations read $\left[-\frac{1}{2}\nabla^{2}+V_{\mathrm{eff}}[\{\phi_{i}^{t-1}\}]+\Delta^{t}\right]\phi^{t}_{i}=\varepsilon^{t}_{i}\phi^{t}_{i}$, with the perturbation function $\Delta(\mathbf{r})=\sum_{\mu}d_{\mu}\omega_{\mu}(\mathbf{r})$, its coefficients $d_{\mu}$ are sampled from a normal distribution with a standard deviation decreasing linearly from $0.102$ to $0.002$ over the SCF iterations and then multiplied with the orbital-free basis functions $\{\omega_{\mu}\}$. In the $\{\eta_{\alpha}\}$ basis representation this amounts to
$\left(\mathbf{F}^{t}+\mathbf{\Delta}^{t}\right)\mathbf{C}^{t}$ $=\mathbf{S}\mathbf{C}^{t}\boldsymbol{\varepsilon}^{t}$ (S.4)
$\Delta^{t}_{\alpha\beta}$ $=\sum_{\mu}d^{t}_{\mu}\int\mathrm{d}^{3}r\,\eta_{\alpha}(\mathbf{r})\eta_{\beta}(\mathbf{r})\omega_{\mu}(\mathbf{r}).$ (S.5)
@@
To obtain density coefficients in the orbital-free ansatz (Eq. 2 in the main text), we employ density fitting. The resolution of identity method by Whitten[69] and Dunlap[70] can be used to fit the orbital-free density coefficients $\mathbf{p}=\{p_{\mu}\}_{\mu\in 1,\dots,N_{\mathrm{OF}}}$ to a Kohn Sham density. Here, the discrepancy of the fit is measured by the residual Hartree energy
$E_{\mathrm{H}}[\rho_{\mathrm{KS}}-\rho_{\mathrm{OF}}]$ $=\int\mathrm{d}\mathbf{r}\int\mathrm{d}\mathbf{r^{\prime}}\frac{(\rho_{\mathrm{KS}}(\mathbf{r})-\rho_{\mathrm{OF}}(\mathbf{r}))(\rho_{\mathrm{KS}}(\mathbf{r^{\prime}})-\rho_{\mathrm{OF}}(\mathbf{r^{\prime}}))}{|\mathbf{r}-\mathbf{r^{\prime}}|}$ (S.6)
$=\mathbf{p}^{\top}\mathbf{\tilde{W}}\mathbf{p}-2\mathbf{p}\,\mathrm{tr}(\mathbf{\tilde{L}}\mathbf{\Gamma})+\sum\limits_{\alpha\beta\gamma\delta}\Gamma_{\alpha\beta}\tilde{D}_{\alpha\beta,\gamma\delta}\Gamma_{\gamma\delta}.$ (S.7)
@@
$(f|g)=\int\mathrm{d}\mathbf{r}\int\mathrm{d}\mathbf{r}^{\prime}\frac{f(\mathbf{r})g(\mathbf{r^{\prime}})}{|\mathbf{r}-\mathbf{r^{\prime}}|}.$ (S.8)
@@
As Eq. S.7 is a quadratic form in $\mathbf{p}$, it can be minimized analytically. But in agreement with M-OFDFT[26], we found that also considering the residual external energy in the minimization yields better fitted external and exchange correlation energies, as using the residual Hartree energy alone only leads to a close fit for this energy. The residual external energy reads
$E_{\mathrm{ext}}[\rho_{\mathrm{KS}}-\rho_{\mathrm{OF}}]$ $=\int\mathrm{d}\mathbf{r}\sum\limits_{(Z_{a},\mathbf{R}_{a})\in\mathcal{M}}\frac{-Z_{a}(\rho_{\mathrm{KS}}(\mathbf{r})-\rho_{\mathrm{OF}}(\mathbf{r}))}{|\mathbf{R}_{a}-\mathbf{r}|}$ (S.9)
$=\mathbf{v}_{\mathrm{ext}}^{\top}\,\mathbf{p}-\mathrm{tr}(\Gamma\mathbf{V}_{\mathrm{ext}}),$ (S.10)
with $[{\mathbf{v}_{\mathrm{ext}}}]_{\mu}=E_{\mathrm{ext}}[\omega_{\mu}]$ and $[{\mathbf{V}_{\mathrm{ext}}}]_{\alpha\beta}=E_{\mathrm{ext}}[\eta_{\alpha}\cdot\eta_{\beta}]$ being the external energies of the individual basis functions and $\mathcal{M}$ denoting the molecule geometry, comprising nuclear positions $\mathbf{R}_{a}$ and corresponding charges $Z_{a}$. The objectives of $E_{\mathrm{ext}}[\rho_{\mathrm{KS}}-\rho_{\mathrm{OF}}]=0$ and minimizing $E_{\mathrm{H}}[\rho_{\mathrm{KS}}-\rho_{\mathrm{OF}}]$ can be combined by differentiating Eq. S.7 with respect to $\mathbf{p}$ and coupling with Eq. S.10, yielding an overdetermined linear system111 The M-OFDFT supplementary[26] suggests minimizing the loss $\mathcal{L}(\mathbf{p})=E_{H}[\rho_{\mathrm{OF}}(\mathbf{p})-\rho_{\mathrm{KS}}]+\left(E_{\mathrm{ext}}[\rho_{\mathrm{OF}}(\mathbf{p})]-E_{\mathrm{ext}}[\rho_{\mathrm{KS}}]\right)^{2},$ (S.11) but the available code optimizes the least squares problem as we do. which is solved by minimizing
$\mathcal{L}(\mathbf{p})=\left\lVert\begin{pmatrix}\mathbf{\tilde{W}}\\ \mathbf{v}_{\mathrm{ext}}^{\top}\end{pmatrix}\mathbf{p}-\begin{pmatrix}\mathrm{tr}(\mathbf{\tilde{L}}\mathbf{\Gamma})\\ \mathrm{tr}(\mathbf{\Gamma}\mathbf{V}_{\mathrm{ext}})\end{pmatrix}\right\lVert^{2}$ (S.12)
@@
$E_{\mathrm{tot}}=T_{\mathrm{S}}+E_{\mathrm{XC}}+E_{\mathrm{H}}+E_{\mathrm{ext}}+E_{\mathrm{nuc}},$ (S.13)
@@
$T_{\mathrm{S}}(\mathbf{C})=\sum\limits_{i}\langle\phi_{i}|\hat{T}|\phi_{i}\rangle=\sum_{\alpha,\beta}\sum_{i}C_{\alpha i}C_{\beta i}\langle\eta_{\alpha}|\hat{T}|\eta_{\beta}\rangle.$ (S.14)
Following prior work [26], we calculate the kinetic energy label for the orbital-free density such that the total energy remains constant, $E_{\mathrm{tot}}(\mathbf{p})=E_{\mathrm{tot}}(\mathbf{C})$, yielding
$T_{\mathrm{S}}(\mathbf{p})=T_{\mathrm{S}}(\mathbf{C})+E_{\mathrm{eff}}(\mathbf{C})-E_{\mathrm{eff}}(\mathbf{p}),\qquad E_{\mathrm{eff}}\coloneq E_{\mathrm{H}}+E_{\mathrm{XC}}+E_{\mathrm{ext}},$ (S.15)
@@
$\frac{\delta T_{\mathrm{S}}[\rho]}{\delta\rho(\mathbf{r})}=-V_{\mathrm{eff}}(\mathbf{r})-\Delta(\mathbf{r})+\mu,$ (S.17)
@@
$\nabla_{p_{\nu}}T_{\mathrm{S}}(\mathbf{p})=\int\frac{\delta T_{\mathrm{S}}[\rho]}{\delta\rho(\mathbf{r})}\omega_{\nu}(\mathbf{r})\mathrm{d}\mathbf{r}=\int\left(-V_{\mathrm{eff}}(\mathbf{r})-\Delta(\mathbf{r})+\mu\right)\omega_{\nu}(\mathbf{r})\mathrm{d}\mathbf{r}.$ (S.18)
The unknown chemical potential $\mu$ can be set to zero, as this only yields a gradient contribution orthogonal to the manifold of normalized densities, which is projected out in density optimization (see section S.6). Instead of obtaining the effective potential function $V_{\mathrm{eff}}(\mathbf{r})$ from the coefficients in the orbital basis, we follow M-OFDFT [26] and directly use the effective potential vector given by
$\mathbf{v}_{\mathrm{eff}}(\mathbf{p})=\nabla_{\mathbf{p}}\left(E_{\mathrm{H}}(\mathbf{p})+E_{\mathrm{XC}}(\mathbf{p})+E_{\mathrm{ext}}(\mathbf{p})\right).$ (S.19)
@@
$\nabla_{\mathbf{p}}T_{\mathrm{S}}(\mathbf{p})=-{\mathbf{v}}_{\mathrm{eff}}-\mathbf{W}\mathbf{d},$ (S.20)
@@
$\mathbf{\tilde{F}}^{t}=\sum_{\tau=1}^{t}\pi_{\tau}^{t}\mathbf{F}^{\tau},\qquad\sum_{\tau=1}^{t}\pi_{\tau}^{t}=1.$ (S.21)
@@
$\mathbf{\tilde{V}}_{\mathrm{eff}}^{t}=\sum_{\tau=1}^{t}\pi_{\tau}^{t}\mathbf{V}^{\tau}_{\mathrm{eff}},$ (S.22)
@@
$\tilde{\mathbf{v}}^{t}_{\mathrm{eff}}=\sum_{\tau=1}^{t}\pi_{\tau}^{t}\mathbf{v}^{\tau}_{\mathrm{eff}}.$ (S.23)
@@
S.2 Model overview
The model input consists of a molecular graph with atomic positions $\{\mathbf{R}_{a}\}$ and atom types $\{Z_{a}\}$ as well as the coefficients $\mathbf{p}$, representing the electron density in terms of atom-centered basis functions, cf. Eq. 2 in the main text. As is customary in local canonicalization, we compute an equivariant local coordinate system for each atom based on the relative position of adjacent non-hydrogen atoms. The basis functions $\{\omega_{\mu}\}$ are a product of a radial function and spherical harmonics. As a consequence, the density coefficients transform via Wigner-D matrices under 3D rotations. To achieve invariance w.r.t. the global orientation of the molecule, the coefficients are transformed into the respective local frame at each atom. As in M-OFDFT[26], the coefficients undergo a “natural reparametrization” into an orthonormal basis, that is specific to the molecule geometry (see section S.3 for details); and the coefficients are rescaled dimensionwise to standardize the model input and the desired gradient range of the model w.r.t. the input coefficients.
@@
$\text{ShrinkGate}(\tilde{\mathbf{p}})=\lambda_{\mathrm{out}}\tanh(\lambda_{\mathrm{in}}\,\tilde{\mathbf{p}}),$ (S.24)
@@
$\tilde{e}_{ij}=\eta_{\mathrm{mul}}(Z_{i},Z_{j})\|\mathbf{r}_{i}-\mathbf{r}_{j}\|+\eta_{\mathrm{bias}}(Z_{i},Z_{j})\,,$ (S.25)
$e_{ij}^{k}=\frac{1}{\sqrt{2\pi}\sigma_{k}}\exp\left(\frac{(\tilde{e}_{ij}-\mu_{k})^{2}}{2\sigma_{k}^{2}}\right).$ (S.26)
@@
Further, an embedding of the atom number $Z$ and an aggregation of edge features over neighboring nodes $\operatorname{MLP}\left(\sum_{j\in\mathcal{N}(i)}e_{ij}^{k}\right)$ are added to the node features. These are then passed to a message-passing graph neural network. At the core of the architecture, we apply 4 (or 8 for QMugs) Graphormer blocks [45]. For experiments with local architectures, we propagate messages along a graph with a radial cutoff of 6 Bohr and otherwise use the fully connected graph. Before each attention block, we also apply a node-wise layer norm. Finally, we employ an energy MLP which produces an atom-specific energy contribution from the final node features. Combined with an atom-specific contribution (“atomic reference module”, see below) based on the statistics of the data, the individual energies of each molecule are aggregated by summation into the total energy prediction. For the QMugs model, we used a hierarchical energy readout (similar to prior work [74]), where, instead of a single energy MLP after the final layer, we employ an energy MLP after every second transformer block to predict atom-specific energy contributions. In the end, all energy contributions of the individual readouts are summed into the final atom-wise energies. Having readout MLPs also in intermediate layers, where the field of view is still small, enforces the prediction of more local energy contributions whereas MLPs after later layers are expected to capture increasingly non-local effects.